### v2

# Machine Learning Detection of Lithium Plating in Lithium-ion Cells: A Gaussian Process Approach

Abstract
Lithium plating during fast charging is a critical degradation mechanism that accelerates capacity fade and can trigger catastrophic safety failures. Recent work has identified a distinctive dQ/dV peak above 4.0 V as a reliable signature of plating onset; however, conventional methods for computing dQ/dV rely on finite differencing with filtering, which amplifies sensor noise and introduces bias in peak location. In this paper, we propose a Gaussian Process (GP) framework for lithium plating detection by directly modeling the charge–voltage relationship Q(V) as a stochastic process with calibrated uncertainty. Leveraging the property that derivatives of GPs remain GPs, we infer dQ/dV analytically and probabilistically from the posterior, enabling robust detection without ad hoc smoothing. The framework provides three key benefits: (i) noise-aware inference with hyperparameters learned from data, (ii) closed-form derivatives with credible intervals for uncertainty quantification, and (iii) scalability to online variants suitable for embedded BMS. Experimental validation on Li-ion coin cells across a range of C-rates (0.2C–1C) and temperatures (0–40 °C) demonstrates that the GP-based method reliably detects plating peaks under low-temperature, high-rate charging, while correctly reporting no peaks in baseline cases. The concurrence of GP-identified differential peaks, reduced charge throughput, and capacity fade measured via reference performance tests confirms the method’s accuracy and robustness, establishing a practical pathway for real-time lithium plating detection.
I INTRODUCTION
Fast charging is pivotal not only for accelerating electric-vehicle adoption but also for electrified aerospace applications—such as satellites—where mission timelines and energy availability demand rapid and reliable charge acceptance [1, 2]. In orbit, repeated short sun–eclipse cycles and limited thermal control expose cells to wide temperature swings, and charging under sub-ambient conditions is a well-established trigger for metallic lithium (Li) plating, a detrimental degradation phenomenon [1, 2, 3, 4]. During high-rate charging, instead of being inserted into the anode as in the nominal reaction, lithium ions can be reduced and deposited as metallic Li on the anode surface. This loss of cyclable lithium accelerates capacity fade, while plated Li may evolve into dendrites that penetrate the separator, creating internal short circuits with catastrophic risks such as fire or explosion. Accordingly, Li plating is recognized as one of the most critical challenges of fast charging, making its early detection and prevention essential for both safety and lifetime [2, 3].
In the electrochemical research community, lithium plating has been extensively studied using a wide range of laboratory-based diagnostic techniques [2]. The most straightforward approach is post-mortem analysis, in which the cell is opened and the metallic lithium deposits are directly observed on the electrode surface [5]. Other in-situ or operando methods such as neutron diffraction, nuclear magnetic resonance (NMR), and X-ray diffraction (XRD) have also been employed to characterize plating and its impact on electrode structure [6, 7, 8]. While powerful, these techniques are either destructive, highly intrusive, or reliant on expensive and complex instrumentation that limits their applicability outside of specialized research laboratories. This has motivated a strong interest in voltage-based diagnostics, which are non-intrusive and thus more suitable for practical battery management in field applications. Most voltage-based approaches reported to date rely on analysis of relaxation or discharge curves, grounded in the theory that plated lithium is oxidized (stripped) first at an approximately constant potential [9, 10, 11]. This process produces a characteristic voltage plateau, which can then be used as a signature of lithium plating. The limitation, however, is that such methods only reveal plating after charging has ended, i.e., during the subsequent rest or discharge phase. As a result, they do not provide true real-time detection of lithium plating onset, which is essential if one aims to actively prevent plating during fast-charging operation.
Recent studies have reported new voltage features that emerge during charging [12, 13]. In particular, differentiating the accumulated charge $Q$ (the integral of current) with respect to voltage $V$ yields a $dQ/dV$ curve that exhibits a distinct peak above a threshold voltage, which has been directly linked to the onset of lithium plating. Operando experiments using novel capillary cell setups have confirmed this finding by directly visualizing the growth of metallic Li deposits [12]. This discovery opens up a promising pathway toward real-time detection and control of lithium plating: in principle, one could compute $dQ/dV$ from routine current and voltage measurements and terminate charging when the peak is observed. However, practical implementation poses significant challenges. The peak in the $dQ/dV$ curve reflects the flattening of the voltage trajectory, since mathematically $dQ/dV$ is the inverse of the time derivative of voltage under constant-current charging. As is well known in control theory, numerical differentiation of signals (e.g., via finite differencing) significantly amplifies sensor noise, leading to large variance in the estimated derivative. While laboratory demonstrations can mitigate this issue with high-precision instrumentation and extensive post-processing (e.g., filtering), robust and accurate detection of $dQ/dV$ peaks based on near-zero voltage derivatives remains extremely difficult in practical settings.
In this work, we propose a Gaussian Process (GP) framework for lithium plating detection by modeling the charge–voltage relationship $Q(V)$ as a GP, directly using the charging voltage curve as input. A key property of GPs is that their derivatives are themselves GPs and remain jointly distributed with the original process [14, 15]. As a result, $dQ/dV$ can be inferred analytically and probabilistically from the GP posterior, without resorting to ad hoc numerical differencing. Compared to the widely adopted filtering–differentiation methods for $dQ/dV$ analysis, such as the Savitzky–Golay (SG) filter [16], the GP approach offers several advantages: (i) GP inference provides a principled treatment of measurement noise, where the hyperparameters (such as the length-scale representing smoothing window and variance representing noise level in data) are self-learned from data, while the traditional filtering approach requires manual tuning of window length and filtering parameters; (ii) GP derivatives are obtained in closed form, thereby avoiding the variance amplification inherent to finite differencing; and (iii) The GP framework yields not only point estimates of $dQ/dV$ but also credible intervals, enabling explicit quantification of uncertainty in peak detection. The probabilistic nature of GPs further makes the method naturally extensible to online detection and adaptive control, where uncertainty-aware decision making is critical. We then build a peak-detection routine with confidence intervals that uses derivative structure and uncertainty bands to distinguish plating-related features from graphite/cathode staging. We demonstrate the method under fast-charge and low-temperature conditions across baseline (no-plating) and plating datasets, showing that GP-based solution improves sensitivity and reliability in the presence of sensor noise.
II Mechanism and Lithium Plating and Electrochemical Features
Lithium plating occurs when, during charging, lithium ions are reduced and deposited as metallic lithium on the graphite anode rather than intercalating into its structure [2]. This parasitic process is favored under conditions of high current or low temperature, where polarization drives the anode potential sufficiently negative relative to Li/Li+. Once initiated, plating accelerates rapidly, consuming cyclable lithium and producing deposits that may evolve into dendrites capable of piercing the separator and triggering internal short circuits. As a result, lithium plating represents a uniquely fast and catastrophic degradation pathway compared with gradual mechanisms such as solid-electrolyte interphase (SEI) growth.
Incremental capacity analysis, or $dQ/dV$ analysis, has long been a standard tool for diagnosing battery degradation [17]. Under low-rate, near-equilibrium charging conditions, the VVQQ curve reflects the thermodynamics of electrode reactions, and its derivative produces a spectrum of peaks corresponding to distinct electrochemical processes such as graphite staging transitions or cathode phase changes. Over aging, systematic peak shifts and broadening provide sensitive markers for gradual degradation mechanisms, including loss of active material (LAM) and loss of lithium inventory (LLI).
Lithium plating, however, differs fundamentally from these slow mechanisms. It progresses rapidly and can cause severe capacity loss within only a few charge–discharge cycles. Failure to detect plating early not only shortens cell lifetime but also introduces serious safety risks. Recent experimental studies have revealed that lithium plating produces a distinctive new peak in the incremental capacity spectrum, typically appearing above $4.0~\mathrm{V}$ across a wide range of cell chemistries [12, 13]. This feature arises from the abrupt change in current partitioning once the anode potential crosses the lithium deposition threshold. Consequently, the presence of a $dQ/dV$ peak above $4.0~\mathrm{V}$ has emerged as a strong electrochemical signature for the onset of plating.
Despite its promise, applying $dQ/dV$ analysis for plating detection under practical fast-charging conditions remains challenging. At high C-rates, overpotentials and IR drop distort the voltage curve, shifting and broadening differential features. More critically, numerical differentiation of noisy voltage–current measurements amplifies sensor noise, producing fluctuations that obscure the plating peak. Traditional filtering methods such as Savitzky-Golay [16, 17], can mitigate variance but introduce their own biases, making robust peak detection under realistic charging conditions an open challenge.
III Gaussian Process-based Lithium Plating Detection Framework
In this section, we describe our GP-based framework for estimating incremental capacity ($dQ/dV$). Incremental capacity curves are obtained as derivatives of discretely sampled voltage–charge data, where the most direct approach is finite differencing. Because numerical differentiation is intrinsically noise-amplifying, it often produces strong fluctuations and errors. To mitigate this, pre-filtering or smoothing is typically applied prior to differencing. However, finite differences effectively act as a high-pass filter, amplifying quantization noise and baseline drift, which leads to spurious ringing and peak shifts even under smoothing [16, 18].
Gaussian Processes (GPs) offer a principled alternative: they model the underlying charge–voltage map as a smooth stochastic function with calibrated uncertainty, and enable noise-robust derivative inference via closed-form conditioning rather than unstable differencing [14, 15, 19]. Concretely, GPs (i) capture nonlinear yet smooth electrochemical responses; (ii) provide credible intervals that quantify confidence in the presence and location of plating-related peaks; and (iii) admit scalable variants (e.g., state-space, sparse, or variational GPs) suitable for online deployment in battery management systems (BMS) for both EVs and spacecraft [20, 21].
III-A Model Formulation: GP over Q(V)
We consider a GP model formalism as
$\begin{split}y&=f(x)+\varepsilon,\;\varepsilon\sim\mathcal{N}(0,\sigma_{n}^{2})\\ f&(x)\sim\mathcal{GP}(m(x),k(x,x^{\prime})).\end{split}$ (1)
Specifically in our case, the input $x$ is the battery voltage $V$, the output $y$ is the accumulated charge $Q$ during charging
$Q=\int_{0}^{t}Idt,$ (2)
which is modeled as a function of $V$ in $f(x)$ with additive zero-mean Gaussian noise $\varepsilon$ with variance $\sigma_{n}^{2}$.
The function $f(x)$ is characterized by a Gaussian process, with mean and variance dependent on $x$. Specifically, it means that any pair of $f(x)$ and $f(x^{\prime})$ are joint Gaussian with respective mean and covariance. Here, the mean is set to zero as a common practice, while the (co-)variance is characterized by a kernel function $k(x,x^{\prime})$. The kernel function encodes the similarity between the input $x$ of interest and other (e.g. training) inputs in the space $x^{\prime}$, thereby determining how information from $f(x^{\prime})$ contributes to the prediction of $f(x)$. In this work, we use the classical squared-exponential kernel,
$k(x,x^{\prime})=\sigma_{f}^{2}\exp\!\left(-\frac{(x-x^{\prime})^{2}}{2\ell^{2}}\right),$ (3)
where the length scale $l$ controls the smoothness of the function, and $\sigma^{2}_{f}$ controls the magnitude of variation among data [14].
The hyperparameters of the model, namely $\theta=\{\ell,\sigma_{f},\sigma_{n}\}$, will be trained by maximizing the log marginal likelihood of the prediction over the training data set, which takes the following form under our GP formulation [14],
$\max_{\theta}-\frac{1}{2}Y^{\top}K_{n}^{-1}(\theta)Y-\frac{1}{2}\log|K_{n}(\theta)|-\frac{N}{2}\log(2\pi).$
In the equation, the vector ${Y}=[y_{1},y_{2},\cdots,y_{N}]$ represents the training output dataset, ${X}=[x_{1},x_{2},\cdots,x_{N}]$ consists of the training input data, and $K_{n}=K({X,X})+\sigma_{n}^{2}$ is the covariance matrix over the training data with each element $K(X,X]_{ij}=k(x_{i},x_{j})$ specified by the kernel in Eqn. (3).
III-B Inference of $dQ/dV$
In our application, the primary objective is to obtain $dQ/dV$, i.e., the derivative of the output with respect to the input,
$f^{\prime}(x)=\frac{dy}{dx}.$
To this end, we leverage a key property of Gaussian Processes as specified in Theorem 1: the derivative of a GP is itself a GP, and it is jointly Gaussian with the original process [14, 15]. This property allows us to infer $dQ/dV$ directly from input–output data without resorting to unstable numerical differencing.
Theorem 1 (Derivative GP and Joint Gaussianity [14]).
Let
$f\sim\mathcal{GP}\!\big(m(\cdot),k(\cdot,\cdot)\big),\quad\mathcal{X}\subset\mathbb{R},$
and assume $f$ is mean-square differentiable (equivalently, $k$ is $C^{2}$ near the diagonal). Then the derivative process
$f^{\prime}(x)=\frac{d}{dx}f(x)$
is also a Gaussian process with mean and covariance
$m^{\prime}(x)=\frac{d}{dx}m(x),\qquad k^{\prime}(x,x^{\prime})=\frac{\partial^{2}}{\partial x\,\partial x^{\prime}}k(x,x^{\prime}).$
Moreover, $f$ and $f^{\prime}$ are jointly Gaussian. For any finite sets
$X=\{x_{1},\dots,x_{n}\},\quad X^{\prime}=\{x^{\prime}_{1},\dots,x^{\prime}_{m}\},$
the random vector
$\begin{bmatrix}f(X)\\[2.0pt] f^{\prime}(X^{\prime})\end{bmatrix}=\begin{bmatrix}f(x_{1}),\dots,f(x_{n}),\,f^{\prime}(x^{\prime}_{1}),\dots,f^{\prime}(x^{\prime}_{m})\end{bmatrix}^{\top}$
is multivariate normal with mean
$\begin{bmatrix}m(X)\\[2.0pt] m^{\prime}(X^{\prime})\end{bmatrix}$
and block covariance
$\begin{bmatrix}K(X,X)&K^{\prime}(X,X^{\prime})\\[4.0pt] K^{\prime}(X^{\prime},X)&K^{\prime\prime}(X^{\prime},X^{\prime})\end{bmatrix}$,
where the blocks are defined element-wise by
$\begin{split}\big[K(X,X)\big]_{ij}&=k(x_{i},x_{j}),\\ \big[K^{\prime}(X,X^{\prime})\big]_{ij}&=\frac{\partial}{\partial x^{\prime}_{j}}k(x_{i},x^{\prime}_{j}),\\ \big[K^{\prime}(X^{\prime},X)\big]_{ij}&=\frac{\partial}{\partial x^{\prime}_{i}}k(x^{\prime}_{i},x_{j}),\\ \big[K^{\prime\prime}(X^{\prime},X^{\prime})\big]_{ij}&=\frac{\partial^{2}}{\partial x^{\prime}_{i}\,\partial x^{\prime}_{j}}k(x^{\prime}_{i},x^{\prime}_{j}).\end{split}$ (4)
The above theorem follows from the fundamental property that any linear operation applied to a Gaussian Process yields another Gaussian Process, and differentiation is a linear operator. Therefore, in our setting we can directly formulate a new GP corresponding to the derivative process, tailored to the estimation of $dQ/dV$,
[…]
where ${Y}=[f(y_{1}),f(y_{2}),\cdots,f(y_{N})]$ are the $Q(V)$ data, and ${F^{\prime}(X^{*})}=[f^{\prime}(x^{*}_{1}),f^{\prime}(x^{*}_{2}),\cdots,f^{\prime}(x^{*}_{M})]$ are the $dQ/dV$ that need to be predicted at the target voltage $V^{*}_{i}$’s. Meanwhile, based on the squared-exponential kernel, the covariance matrices are given as
$\begin{split}[K^{\prime}({X,X^{*}})]_{ij}&=\partial_{x^{*}}k(x_{i},x^{*}_{j})=k(x_{i},x^{*}_{j})\frac{(x_{i}-x^{*}_{j})}{\ell^{2}}\\ [K^{\prime\prime}(X^{*},X^{*})]_{ij}&=\partial_{x}\partial_{x^{*}}k(x_{i},x^{*}_{j})\\ &=k(x_{i},x^{*}_{j})\left(\frac{1}{\ell^{2}}-\frac{(x_{i}-x^{*}_{j})^{2}}{\ell^{4}}\right).\end{split}$ (6)
Finally, based on the new joint Gaussian process, we can provide a posterior estimate of the $dQ/dV$ curve $F^{\prime}(X^{*})$} conditional on the noisy $\mathbf{y}$ data as
$F^{\prime}(x^{*})|Y\sim\mathcal{N}(\mu^{\prime*},\Sigma^{\prime*})$
with
$\mu^{\prime}(X^{*})$ $=K^{\prime}({X^{*},X})K(X,X)^{-1}Y,$
$\Sigma^{\prime*}$ $=K^{\prime\prime}({X^{*},X^{*}})-K^{\prime}({X_{*},X})K(X,X)^{-1}K^{\prime}({X,X_{*}}).$
where, the mean $\mu^{\prime}(X^{*})$ gives the estimate of the $dQ/dV$, while $\Sigma^{\prime*}$ can be used to construct the confidence interval of the estimate.
IV Experimental Validation
IV-A Testing Setup
Figure 1: Representative test sequence (1C at $25\,^{\circ}\mathrm{C}$) showing the complete protocol used for experiments. Degradation cycles consisting of repeated constant-current charge/discharge cycles (left), followed by a Reference Performance Test (RPT) consisting of a low-rate capacity test and Hybrid Pulse Power Characterization (HPPC) pulses (right). Blue = voltage (left axis); red = current (right axis).
To validate the proposed lithium plating detection scheme, we conducted constant-current ($C$) charge/discharge aging cycles under multiple operating conditions. Charging was performed at C-rates of 1C, 0.8C, 0.6C, 0.4C, and 0.2C, across ambient temperatures ranging from $40\,^{\circ}\text{C}$ down to $0\,^{\circ}\text{C}$ (inclusive). Each charging cycle was followed by discharging at 1C down to the manufacturer-specified lower cutoff voltage $V_{\min}=2.75~\text{V}$, with an upper cutoff voltage $V_{\max}=4.2~\text{V}$. A rest period of $5~\text{min}$ was inserted at the end of each cycle, and at least 10 cycles were repeated under each condition. The experiments were performed using an Arbin LBT 21084 battery cycler for charge/discharge control and data acquisition, together with an ESPEC environmental chamber for temperature regulation. The tested cell was a 45 mAh LIR 2032 Li-ion coin cell. To track degradation, periodic Reference Performance Tests (RPTs) were conducted, including 1/10 C constant-current capacity tests and hybrid pulse power characterization (HPPC) to evaluate internal resistance. A typical testing sequence is shown in Fig. 1.
IV-B Results and Discussion
Figure 2: GP-based plating detection across charge rates and temperatures. (a) V(t) during $C$ charge for representative cycles (legend shows C-rate and temperature). (b) capacity–voltage trajectories Q(V). (c) GP-derived incremental capacity dQ/dV with inset figures (i), (ii) and (iii) showing the magnified 95% credible bands at the start of the cycle with bounds as % of peak magnitude of the no-plating and plating clusters, respectively.
Figure 3: Normalized Charge Throughput values for each $C$ (constant-current) charge cycle across charging rates and temperatures for the first 10 cycles
Using voltage–current data collected during $C$ charging, we trained GP models on the measured $Q(V)$ trajectory and inferred $dQ/dV$ directly from the derivative GP posterior. This probabilistic differentiator yields smooth incremental-capacity curves with calibrated uncertainty, enabling robust peak finding under realistic sensor noise. Fig. 2 illustrates the full GP-based analysis workflow for several charge-rate/temperature combinations. The first subplot shows $V(t)$ during $C$ charge, where increasing C-rate and decreasing temperature shorten charge time due to increased polarization. The second subplot maps the data to $Q(V)$ and overlays the GP posterior. The third subplot presents the GP-derived $dQ/dV$ (mean and $\pm$95% credible interval), revealing transient differential peaks that are otherwise obscured by finite-difference noise.
TABLE I: Summary of operating conditions, GP-based differential peak detection, plating classification, charge throughput, and capacity degradation rates.
Operating
condition
Peak
>> 4.0 V?
Lithium
plating?
Charge throughput
decrease rate
(% loss/cycle)
Capacity
decrease rate
(% loss/cycle)
(from RPT)
$1.0\,\mathrm{C}$ @ $10^{\circ}\mathrm{C}$ Yes Yes 1.464 0.877
$0.8\,\mathrm{C}$ @ $10^{\circ}\mathrm{C}$ Yes Yes 2.734 1.979
$0.6\,\mathrm{C}$ @ $10^{\circ}\mathrm{C}$ Yes Yes 2.038 1.692
$0.4\,\mathrm{C}$ @ $10^{\circ}\mathrm{C}$ Yes Yes 1.671 1.237
$0.6\,\mathrm{C}$ @ $0^{\circ}\mathrm{C}$ Yes Yes 3.617 1.175
$0.4\,\mathrm{C}$ @ $0^{\circ}\mathrm{C}$ Yes Yes 1.711 0.845
$0.2\,\mathrm{C}$ @ $0^{\circ}\mathrm{C}$ No No 0.029 0.094
$1.0\,\mathrm{C}$ @ $40^{\circ}\mathrm{C}$ No No 0.093 0.016
$1.0\,\mathrm{C}$ @ $25^{\circ}\mathrm{C}$ No No 0.227 0.031
Fig. 2 shows that, across the tested operating conditions, the GP-based $dQ/dV$ analysis reveals a clear separation into two clusters: one cluster exhibits a distinct peak above $4.0~\mathrm{V}$, while the other shows no such peak. According to prior studies [12], a peak above $4.0~\mathrm{V}$ is the characteristic electrochemical signature of lithium plating. Closer inspection confirms that all cases with such peaks correspond to high charging rates under low-temperature conditions—specifically, six cases with charging rates above $0.4C$ and temperatures at or below $10\,^{\circ}\text{C}$. In contrast, the cluster without a peak above $4.0~\mathrm{V}$ comprises three cases under moderate charging or temperature conditions: $0.2C$ at $0\,^{\circ}\text{C}$, and $1C$ at $25\,^{\circ}\text{C}$ or $40\,^{\circ}\text{C}$.
To validate these observations, we further examined the evolution of capacity across cycles under each condition using two metrics. First, we considered the charge throughput of each charging cycle, i.e., the total energy charged into the battery per cycle, as shown in Fig. 3. All six cases highlighted in the bottom cluster in Fig. 3 showed significant and continuous decrease in charge throughput, with losses of at least 13% over just 10 cycles (i.e., degradation rates exceeding 1.3% per cycle). The most severe case was $0.6C$ in $0\,^{\circ}\text{C}$, which lost nearly 30% throughput after 10 cycles. Such abnormal loss rates are consistent with lithium plating, where cyclable lithium is rapidly consumed. By contrast, the three cases highlighted in the second cluster (top) exhibited stable throughput: after a brief initial stabilization, throughput remained essentially constant over repeated cycles, with no indication of plating. We have also performed periodic Reference Performance Tests (RPTs) to measure exact capacity at $C/10$, which corroborated the throughput analysis: conditions associated with plating showed marked capacity fade, while non-plating conditions retained near-constant capacity over cycles.
Table I summarizes all the findings. In plating cases, charge throughput and capacity loss (as a %/cycle) are higher by about an order of magnitude compared with non-plating cases. Our GP-based method successfully identified peaks above $4.0~\mathrm{V}$ in all cases where plating occurred, while correctly reporting no peaks under non-plating conditions. Taken together, (i) the presence of $dQ/dV$ peaks above $4.0~\mathrm{V}$, (ii) the observed reduction in charge throughput over aging cycles, and (iii) the increased capacity fade confirmed by RPT measurements validate the accurate detection capability of our method.
Finally, it is important to note the role of GP uncertainty quantification. Across all plating cases, the credible intervals around the detected peaks were exceptionally tight as shown in the insets of Fig. 2(c). On average, the bounds were within 0.3% of the peak magnitude, indicating very high confidence in the detection. This level of certainty is unattainable with conventional methods, which provide only smoothed (point) estimates without any measure of confidence.
V Conclusions
We have presented a voltage-only method for lithium plating detection that employs an exact Gaussian Process (GP) prior for capacity as a function of voltage, $Q(V)$. By analytically computing $dQ/dV$ from the joint value–derivative posterior, the method avoids the noise amplification inherent to finite-difference ICA/DVA and additionally provides calibrated uncertainty as a reliability metric.
Across all tested conditions, the GP-based analysis yields a clear separation between plating and non-plating cases based on the location of the dominant incremental-capacity peak. Cycles whose $dQ/dV$ maximum occurs below $4.0~\mathrm{V}$ are consistently plating-negative, whereas cycles with a dominant peak above $4.0~\mathrm{V}$ near end-of-charge show plating. Baseline conditions ($1\mathrm{C}$ at $25^{\circ}\mathrm{C}$ and $40^{\circ}\mathrm{C}$) exhibit no high-confidence peaks above $4.0~\mathrm{V}$, while colder and/or higher-rate conditions produce pronounced $4.0~\mathrm{V}$ peaks with tight uncertainty bands, consistent with plating.
These differential signatures align closely with independent performance metrics. Plating-positive cases exhibit faster declines in delivered charge throughput per $C$ cycle as well as accelerated capacity fade confirmed by periodic RPTs, with the coldest and highest-rate cases degrading most severely. In contrast, baseline and non-plating cases show stable charge throughput and much slower capacity fade.
Taken together, the concurrence of (i) high-confidence $dQ/dV$ peaks localized above $4.0~\mathrm{V}$, (ii) reduced charge throughput over cycling, and (iii) accelerated capacity fade provides a coherent, physics-consistent picture linking the GP-identified signature to practically relevant performance loss. These results demonstrate that exact GP–based voltage analysis enables more robust lithium plating diagnostics, paving the way for reliable online detection and integration into advanced battery management systems.
ACKNOWLEDGMENT
We appreciate the funding support from the USAF Proteus Space STTR and the NSF CAREER Program (Grant No.2046292).